OpenStudy (anonymous):

the position of a moving particle is given s(t)=2t^3-21t^2+60t where distance is measured in ft & time in secs. find velocity of particle at t=5

OpenStudy (anonymous):

v=ds/dt=6t^2-42t+60 now put t=5, it will be.6.(25)-42.(5)+60=0 ans =0